OpenStudy (anonymous):

I am having trouble figuring out how to differentiate x/2. Please help!

OpenStudy (welshfella):

Write it as 0.5x Then use the formula for algebraic differentiation

OpenStudy (anonymous):

ok, how do you you get it to 0.5x ? Which exponent rule is that?

OpenStudy (welshfella):

F(x) = ax^n F ' x = an x ^(n-1)

OpenStudy (anonymous):

Oh wait, I dont think exponent rules have anything to do with it. I just meant what made that possible

OpenStudy (welshfella):

X/2 = 1/2 x = 0.5 x If you prefer fractions use 1/2 x

OpenStudy (welshfella):

I prefer decimals

OpenStudy (anonymous):

Ah ok. Thanks!

OpenStudy (welshfella):

So here w e have a=0.5, n=1 and of course n-1 = 0

OpenStudy (welshfella):

Yw